a,b,c are three distinct real numbers and there are real numbers x,y such that a^3+ax+y=0,b^3+bx=y=0 and c^3+cx+y=0 Show that a+b+c=0
Answer by Edwin McCravy(20054)

a,b,c are three distinct real numbers and there are real numbers x,y such that a3+ax+y=0,b3+bx=y=0 and c3+cx+y=0 Show that a+b+c=0

All three equations are cases of the cubic 
polynomial equation:

z3+xz+y=0

Since a3+ax+y=0,b3+bx=y=0 and c3+cx+y=0, this shows
that z=a, z=b and z=c are distinct solutions to 

z3+xz+y=0

and since that equation is of degree 3 it can have no
more than 3 solutions. Thus z=a, z=b, and z=c are the 
three real solutions to

z3+xz+y=0

Therefore its left side must factor as

(z-a)(z-b)(z-c)=0

(z-a)(z2-cz-bz+bc)=0

z3-cz2-bz2+bcz-az2+acz+abz-abc=0

z3-(a+b+c)z3+(ab+ac+bc)z-abc=0

This must be the same equation as

z3+xz+y=0

Therefore the coefficient of z2 in

z3-(a+b+c)z2+(ab+ac+bc)z-abc=0

must be 0 since z3+xz+y=0 has no
term in z2.  Therefore

-(a+b+c)=0 or

a+b+c=0.

That's all you were asked to show.

However by equating the other terms,
you can also find x and y:

x=ab+ac+bc and y=-abc.

Edwin